OpenStudy (anonymous):

evaluate the expression log3(1/27)

OpenStudy (anonymous):

\[\log_{3}(\frac{ 1 }{ 27 }) \]

OpenStudy (anonymous):

help please!

OpenStudy (kropot72):

\[\large \log_{3} \frac{1}{27}=x\] Then by the rules of logs: \[\large 3^{x}=\frac{1}{27}=\frac{1}{3^{3}}=3^{-3}\] Can you now see the value of x.

OpenStudy (kropot72):

\[\large 3^{x}=3^{-3}\] So what must the value of x be?

OpenStudy (kropot72):

For example \[\large 10^{x}=10^{7}\] therefore x = 7.

OpenStudy (kropot72):

So if \[\large 3^{x}=3^{-3}\] what will be the value of x?

OpenStudy (kropot72):

@radicalgal Are you there?

OpenStudy (anonymous):

yes sorry.. so then the -3 is the answer

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

thank you!

OpenStudy (kropot72):

You're welcome :)